Jet Corporation had 8 million shares of common stock outstanding during the current calendar year. On July 1, Jet issued ten thousand $1,000 face value, convertible bonds. Each bond is convertible into 50 shares of common stock. The bonds were issued at face amount and pay interest semiannuallly for 20 years. They have a stated rate of 12%. Jet had net income of $18 million (after a 40% tax). Jet would report the following EPS data (rounded):

Basic EPS                           Diluted EPS

  1. $2.25                                   $2.23
  2. $2.23                                   $2.23
  3. $2.25                                   $2.25
  4. $2.25                                   $2.16
  5. $2.23                                   $2.16
Answer #1

Answer:2.25,2.23

Basic EPS=18,000,000/8000,000=2.25

Diluted EPS=(18,000,000+(600,000*(1-40%))/8000,000+(10000*50*6/12)=2.23